OpenStudy (anonymous):

f(x) = sqrt( 3x + 7) , g(x) = sqrt(3x - 7 ) Find (f + g)(x)

OpenStudy (e.mccormick):

Where are you getting stuck on this?

OpenStudy (anonymous):

The whole thing hahah (:

OpenStudy (e.mccormick):

Do you know what the \((f + g)(x)\) part means?

OpenStudy (anonymous):

adding

OpenStudy (e.mccormick):

Yes, but, I meant more like this: \((f + g)(x) = [f(x)] + [g(x)]\) \((f - g)(x) = [f(x)] - [g(x)]\) \((f \times g)(x) = [f(x)] \times [g(x)]\) \((f \div g)(x) = [f(x)] \div [g(x)]\) AKA: \(\left( \dfrac{f}{g}\right) (x) = \dfrac{[f(x)]}{[g(x)]}\) \((f \circ g)(x) = f([g(x)]) \)

OpenStudy (anonymous):

Yes I know that

OpenStudy (e.mccormick):

OK. So what does it become when you write it out as addition?

OpenStudy (anonymous):

sqrt(6x) ?

OpenStudy (anonymous):

^ is this the answer

OpenStudy (e.mccormick):

Rememebr your rules for adding roots?

OpenStudy (e.mccormick):

The parts under the radical must match or they cannot be added.

OpenStudy (anonymous):

okay

OpenStudy (e.mccormick):

And that makes it?

OpenStudy (anonymous):

sqrt(6x) ?

OpenStudy (e.mccormick):

It is not \(\sqrt{6x}\) because \(( 3x + 7) \ne (3x - 7)\) they can not be added into one thing under a radical.

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

do I just leave it ?

OpenStudy (e.mccormick):

Yah... they snuck in a simple one. =)

OpenStudy (e.mccormick):

\(\sqrt{3x + 7} + \sqrt{3x - 7 }\)

OpenStudy (e.mccormick):

If it was times between them, it would be a whole different thing! But for \(\pm\) it just sits there because they do not match.

OpenStudy (anonymous):

Aww thanks (:

OpenStudy (e.mccormick):

np. Have fun!